OpenStudy (anonymous):

@Nnesha Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1 fourthx + 2. y = 4x - 1 y = 4x - 3 y = 1 fourthx - 3 y = 1 fourthx - 1

OpenStudy (anonymous):

my firmed helped em out on this one she said it was a is she right ?

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -negative one halfx + 4. y = 2x + 8 y = 2x y = -2x - 8 y = -2x

OpenStudy (anonymous):

she also said this wa

OpenStudy (anonymous):

c

Nnesha (nnesha):

there is key word " parallel"

Nnesha (nnesha):

which mean parallel lines have same slope do you know the slope of this equation y = 1 fourthx + 2. ??

OpenStudy (anonymous):

@Nnesha may you help me check my home work

Nnesha (nnesha):

yep i said that is not right what is slope of y = 1/4x + 2 ??

Nnesha (nnesha):

y = mx + b slope intercept form equation where m is slope which is ALWAYS with x